# Alternating and symmetric separability of free products

### 19 Apr 2026 v1

@@ Abstract.
Let $F\ast G$ be a free product of a free group $F$ and a LERF group $G$. In this note, we provide sufficient conditions for a subgroup $H$ of $F\ast G$ to be $\mathcal{A}\cup\mathcal{S}$-separable, that is, for any finite set $\{\gamma_{1},\ldots,\gamma_{n}\}\subset(F\ast G)\setminus H$, there is a surjection $f$ from $F\ast G$ to an alternating or symmetric group such that $f(\gamma_{i})\notin f(H)$ for all ii. As a corollary, any finitely generated infinite-index subgroup of a free group is $\mathcal{A}\cup\mathcal{S}$-separable in the free product of the free group and an arbitrary LERF group, generalizing a result of Wilton.
@@ 1. Introduction
Definition $1.$
Let $H$ be a subgroup of a finitely generated group $G$ and let $\mathcal{C}$ be a class of groups. We say that $H$ is $\mathcal{C}$-separable in $G$ if for any finite set $\{\gamma_{1},\ldots,\gamma_{n}\}\subset G\setminus H$, there is a surjective homomorphism $f$ from $G$ to a group in $\mathcal{C}$ such that $f(\gamma_{i})\notin f(H)$ for all ii.
@@
[18, Theorem A] Let $H$ be a finitely generated subgroup of infinite index in the free group $F_{r}$ and let $\{\gamma_{1},\ldots,\gamma_{n}\}$ be any finite set of $F\setminus H$. Then there is a surjection $f$ from $F_{r}$ to some finite alternating group $A_{k}$ such that $f(\gamma_{i})\notin f(H)$ for all ii.
@@ 2.1. Labeled graph and embedding
$2.$ Relative Cayley graph
For a group $G=\langle X\mid R\rangle$, we denote the Cayley graph of $G$ with respect to the given generating set $X$ by $\mathrm{Cay}(G)$. Recall that $\mathrm{Cay}(G)$ is an oriented graph with the vertex set $\mathrm{V}(\mathrm{Cay}(G))=G$ and the edge set $\mathrm{\mathrm{E}}(\mathrm{Cay}(G))=G\times X^{\ast}$. An edge $(g,x)\in\mathrm{E}(\mathrm{Cay}(G))$ originates at $g$ and terminates at gxgx.
@@
and the edge set $\mathrm{E}(\mathrm{Cay}(G,H))=\mathrm{V}(\mathrm{Cay}(G,H))\times X^{\ast}$. We use $\mathrm{Cay}(G,H,H\cdot 1)$ when the base point $H\cdot 1$ is emphasized. An edge $(Hg,x)$ originates from the vertex HgHg and ends at HgxHgx. Moreover, a path $p$ originating from HgHg must end at $Hg\mathrm{Lab}(p)$. Therefore, we have the following.
Lemma $2.$
@@ 2.3. Precover
    $\Gamma$ is called a precover of AA, if it is AA-based, and each $X$-component of $\Gamma$ is a cover of $G_{1}$ and each $Y$-component of $\Gamma$ is a cover of $G_{2}$.
There is an equivalent description of the AA-based graph as follows.
@@
If $A=G_{1}\ast G_{2}$ is a free product, then any graph labeled with $X\cup Y$ is AA-based if and only if each $G_{i}$-component is $G_{i}$-based for $i=1,2$. In particular, it is AA-based if each monochromatic component is a cover of the corresponding free factor.
@@
Let $\Gamma$ be a finite, connected, AA-based graph labeled with $X\cup Y$ and $\Gamma_{1}$ be an $X$-component. Let $P$ be a loop in $\Gamma$ starting from $v_{1}\in\Gamma_{1}$. If $1\neq\mathrm{Lab}(P)\in G_{1}$, then there is a loop $P^{\prime}$ in $\Gamma_{1}$ starting from $v_{1}$ such that $\mathrm{Lab}(P^{\prime})=\mathrm{Lab}(P)$.
@@
and AA is a free product of $G_{1}$ and $G_{2}$, we have $\mathrm{Lab}(P_{i})=1$ for each $P_{i}$ labeled with $Y$. Then $P_{i}$ must be closed because $\Gamma$ is AA-based. Moreover, as $P_{i}$ is maximal, both $P_{i-1}$ and $P_{i+1}$ are contained in $X$-components with
@@ 3.1. Jordan’s theorem
For each $k>2$ there exists NN such that if $n>N$, $G\leq S_{n}$ is a primitive subgroup and there exists $\gamma\in G\setminus\{e\}$, which moves less than $k$ elements, then $G=S_{n}$ or $A_{n}$.
@@
For each $k>2$, there exists NN such that for any prime $p>N$ and any transitive subgroup $G\leq S_{p}$ if there exists $\gamma\in G\setminus\{e\}$, which moves less than $k$ elements, then $G=S_{p}$ or $A_{p}$.
@@ 3.2. Finite cover and Wilton’s construction
We begin by constructing two particular labeled graphs $W_{n}$ and $V_{s}$ with n=pΓkΓ4n=p-k-4 and 44 vertices respectively, such that $x_{i}\in X$ acts trivially on $\mathrm{V}(W_{n})$ for each $i\geq 2$ and $x_{2}$ acts nontrivially on $\mathrm{V}(V_{s})$.
@@ 5. Proof of the main theorem
The last “=" follows from Lemma $2.$ Since $\Gamma$ is the subgraph of $\mathrm{Cay}(F_{r}\ast G,H)$, each loop in $\mathrm{Lab}(\Gamma,H\cdot 1)$ is again a loop in $\mathrm{Lab}(\mathrm{Cay}(F_{r}\ast G,H),H\cdot 1)$, that is,

### 21 Apr 2026 v2

@@ Abstract.
Let $F\ast G$ be a free product of a free group $F$ and a LERF group $G$. In this note, we provide sufficient conditions for a subgroup $H$ of $F\ast G$ to be $\mathcal{A}\cup\mathcal{S}$-separable, that is, for any finite set $\{\gamma_{1},\ldots,\gamma_{n}\}\subset(F\ast G)\setminus H$, there is a surjection $f$ from $F\ast G$ to an alternating or symmetric group such that $f(\gamma_{i})\notin f(H)$ for all $i$. As a corollary, any finitely generated infinite-index subgroup of a free group is $\mathcal{A}\cup\mathcal{S}$-separable in the free product of the free group and an arbitrary LERF group, generalizing a result of Wilton.
@@ 1. Introduction
Definition 1.1.
Let $H$ be a subgroup of a finitely generated group $G$ and let $\mathcal{C}$ be a class of groups. We say that $H$ is $\mathcal{C}$-separable in $G$ if for any finite set $\{\gamma_{1},\ldots,\gamma_{n}\}\subset G\setminus H$, there is a surjective homomorphism $f$ from $G$ to a group in $\mathcal{C}$ such that $f(\gamma_{i})\notin f(H)$ for all $i$.
@@
[18, Theorem A] Let $H$ be a finitely generated subgroup of infinite index in the free group $F_{r}$ and let $\{\gamma_{1},\ldots,\gamma_{n}\}$ be any finite set of $F\setminus H$. Then there is a surjection $f$ from $F_{r}$ to some finite alternating group $A_{k}$ such that $f(\gamma_{i})\notin f(H)$ for all $i$.
@@ 2.1. Labeled graph and embedding
2.2. Relative Cayley graph
For a group $G=\langle X\mid R\rangle$, we denote the Cayley graph of $G$ with respect to the given generating set $X$ by $\mathrm{Cay}(G)$. Recall that $\mathrm{Cay}(G)$ is an oriented graph with the vertex set $\mathrm{V}(\mathrm{Cay}(G))=G$ and the edge set $\mathrm{\mathrm{E}}(\mathrm{Cay}(G))=G\times X^{\ast}$. An edge $(g,x)\in\mathrm{E}(\mathrm{Cay}(G))$ originates at $g$ and terminates at $gx$.
@@ 2.2. Relative Cayley graph
and the edge set $\mathrm{E}(\mathrm{Cay}(G,H))=\mathrm{V}(\mathrm{Cay}(G,H))\times X^{\ast}$. We use $\mathrm{Cay}(G,H,H\cdot 1)$ when the base point $H\cdot 1$ is emphasized. An edge $(Hg,x)$ originates from the vertex $Hg$ and ends at $Hgx$. Moreover, a path $p$ originating from $Hg$ must end at $Hg\mathrm{Lab}(p)$. Therefore, we have the following.
Lemma 2.2.
@@ 2.3. Precover
    $\Gamma$ is called a precover of $A$, if it is $A$-based, and each $X$-component of $\Gamma$ is a cover of $G_{1}$ and each $Y$-component of $\Gamma$ is a cover of $G_{2}$.
There is an equivalent description of the $A$-based graph as follows.
@@
If $A=G_{1}\ast G_{2}$ is a free product, then any graph labeled with $X\cup Y$ is $A$-based if and only if each $G_{i}$-component is $G_{i}$-based for $i=1,2$. In particular, it is $A$-based if each monochromatic component is a cover of the corresponding free factor.
@@
Let $\Gamma$ be a finite, connected, $A$-based graph labeled with $X\cup Y$ and $\Gamma_{1}$ be an $X$-component. Let $P$ be a loop in $\Gamma$ starting from $v_{1}\in\Gamma_{1}$. If $1\neq\mathrm{Lab}(P)\in G_{1}$, then there is a loop $P^{\prime}$ in $\Gamma_{1}$ starting from $v_{1}$ such that $\mathrm{Lab}(P^{\prime})=\mathrm{Lab}(P)$.
@@
and $A$ is a free product of $G_{1}$ and $G_{2}$, we have $\mathrm{Lab}(P_{i})=1$ for each $P_{i}$ labeled with $Y$. Then $P_{i}$ must be closed because $\Gamma$ is $A$-based. Moreover, as $P_{i}$ is maximal, both $P_{i-1}$ and $P_{i+1}$ are contained in $X$-components with
@@ 3.1. Jordan’s theorem
For each $k>2$ there exists $N$ such that if $n>N$, $G\leq S_{n}$ is a primitive subgroup and there exists $\gamma\in G\setminus\{e\}$, which moves less than $k$ elements, then $G=S_{n}$ or $A_{n}$.
@@
For each $k>2$, there exists $N$ such that for any prime $p>N$ and any transitive subgroup $G\leq S_{p}$ if there exists $\gamma\in G\setminus\{e\}$, which moves less than $k$ elements, then $G=S_{p}$ or $A_{p}$.
@@ 3.2. Finite cover and Wilton’s construction
We begin by constructing two particular labeled graphs $W_{n}$ and $V_{s}$ with n=pΓkΓ4n=p-k-4 and $4$ vertices respectively, such that $x_{i}\in X$ acts trivially on $\mathrm{V}(W_{n})$ for each $i\geq 2$ and $x_{2}$ acts nontrivially on $\mathrm{V}(V_{s})$.
@@ 5. Proof of the main theorem
The last “=" follows from Lemma 2.2. Since $\Gamma$ is the subgraph of $\mathrm{Cay}(F_{r}\ast G,H)$, each loop in $\mathrm{Lab}(\Gamma,H\cdot 1)$ is again a loop in $\mathrm{Lab}(\mathrm{Cay}(F_{r}\ast G,H),H\cdot 1)$, that is,